OpenStudy (anonymous):

evaluate the logarithmic function. /log(log10^{10,000})/

OpenStudy (anonymous):

\[\log_{10}(10,000)=5\]

OpenStudy (anonymous):

is this supposed to be \[\log(\log(10^{100000}))\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so then \[\log(10^{100000})=100000\] since your base is ten

OpenStudy (anonymous):

and \[\log(10,000)=5\] is your answer

OpenStudy (anonymous):

no the answer is 4

OpenStudy (anonymous):

at least that's what the book says

OpenStudy (anonymous):

i got it

OpenStudy (anonymous):

\[\log_{10}x=10,000\log_{10}10 \] =10,000 \[\log_{10}10,000=4 \]

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

yes, i miscounted the zeros!